# GaussianOcc: Fully Self-supervised and Efficient 3D Occupancy Estimation with Gaussian Splatting

### v3

@@ Abstract
We introduce GaussianOcc, a systematic method that investigates two usages of Gaussian splatting for fully self-supervised and efficient 3D occupancy estimation in surround views. First, traditional methods for self-supervised 3D occupancy estimation still require ground truth 6D poses from sensors during training. To address this limitation, we propose Gaussian Splatting for Projection (GSP) module to provide accurate scale information for fully self-supervised training from adjacent view projection. Additionally, existing methods rely on volume rendering for final 3D voxel representation learning using 2D signals (depth maps and semantic maps), which is time-consuming and less effective. We propose Gaussian Splatting from Voxel space (GSV) to leverage the fast rendering properties of Gaussian splatting. As a result, the proposed GaussianOcc method enables fully self-supervised (no ground-truth pose) 3D occupancy estimation in competitive performance with low computational cost (2.7 times faster in training and 5 times faster in rendering). The relevant code is available in https://github.com/GANWANSHUI/GaussianOcc.git.
@@
In this supplementary material, we provide more implementation details, experiment results, and further discussion on the limitations and future work.
@@ 1 Introduction
Figure 1: Problem setting of GaussianOcc. Given a surround image sequence and its corresponding 2D semantic annotation, GaussianOcc is able to perform 3D occupancy estimation without the need for ground truth occupancy label and ground truth 6D pose for training.
In recent years, both vision-centric and grid-centric perception approaches have gained significant attention in the autonomous driving industry and academia [34, 39]. Among these, surround view 3D occupancy estimation [31, 45, 51, 48] has emerged as a core perception task and a promising alternative to bird’s-eye view (BEV) methods [9, 28, 29]. To facilitate 3D occupancy estimation, several benchmarks have been developed for supervised training [45, 43, 41, 42], though these require substantial effort in 3D annotation. To reduce the burden of annotation, self-supervised [11, 54, 23, 5, 17] and weakly-supervised [37] learning approaches based on volume rendering have been proposed [35, 46, 10]. Volume rendering allows 3D representation learning using 2D supervision signals, such as 2D semantic maps and depth maps, thereby eliminating the need for extensive 3D annotation during training.
Existing methods [54, 23] achieve self-supervised learning through volume rendering, where the 2D semantic map supervision is derived from open-vocabulary semantic segmentation [55], and the depth map supervision is obtained from self-supervised depth estimation [12]. However, these approaches face two significant limitations. First, volume rendering is performed at real-world scale, which requires the availability of ground truth 6D pose to calculate the multi-view photometric loss across sequential images. Second, the inefficiency in volume rendering poses a challenge, the same as in novel view synthesis tasks [26, 1, 18], due to the dense sampling operation required. These limitations impede the development of a more general and efficient paradigm for self-supervised 3D occupancy estimation.
To address the aforementioned limitations, we propose a fully self-supervised and efficient approach to 3D occupancy estimation using Gaussian splatting [26, 1]. Specifically, we introduce the use of Gaussian splatting to perform cross-view splatting, where the rendered image constructs a cross-view loss that provides scale information during joint training with the 6D pose network. This eliminates the need for ground truth 6D pose during training. To improve rendering efficiency, we move away from the dense sampling required in traditional volume rendering. Instead, we propose performing Gaussian splatting directly from the 3D voxel space. In this approach, each vertex in the voxel grid is treated as a 3D Gaussian, and we optimize the attributes of these Gaussians—such as semantic features and opacity—directly within the voxel space. Through this novel approach, our proposed method, GaussianOcc, makes a progress toward fully self-supervised and efficient 3D occupancy estimation, as outlined in Figure 1.
@@
  • We propose Gaussian splatting for cross-view projection module, which can provide scale information to get rid of the need of ground truth 6D pose during the training.
@@ 2.2 Surround view 3D occupancy estimation
Figure 2: GaussianOcc is a two-stage method. In Stage 1, we train a scale-aware 6D pose network, using a U-Net architecture to predict Gaussian attributes in the 2D image grid space for cross-view Gaussian splatting. This approach provides scale information in the joint training with the 6D pose network. Based on the 6D pose from Stage 1, we perform self-supervised 3D occupancy estimation in Stage 2, where we lift the 2D features to a 3D voxel space and propose voxel grid Gaussian splatting for fast rendering. Note that, for clarity, we omit the line from the 6D pose network to the loss in Stage 1, and the 2D encoder is independent for each stage (not shared).
@@
Note that, two recent works, GaussianFormer [24] and GaussianBeV [6], are closely related to ours in their focus on 3D occupancy estimation and BEV prediction. However, our exploration diverges by focusing on two new properties that Gaussian splatting can contribute to occupancy estimation: scale-aware training and faster rendering.
@@ 3.1 Preliminaries
3D Gaussian splatting [26] is for modeling static 3D scenes using point primitives, where each primitive is characterized with the following attributes: (1) a 3D position $\mathcal{X}\in\mathbb{R}^{3}$, (2) a color defined by SH coefficients $c\in\mathbb{R}^{k}$ (where $k$ denotes the dimensionality of the SH basis), (3) a rotation represented by a quaternion $r\in\mathbb{R}^{4}$, (4) a scaling factor $s\in\mathbb{R}^{3}_{+}$, and (5) an opacity $\alpha\in[0,1]$. The original Gaussian splatting is for scene-specific, fast 3D novel view synthesis, where the attributes of Gaussian points are optimized by the multi-view constraint. In contrast, in this work, we explore a setting where Gaussian attributes are well-aligned in both 2D and 3D grids, allowing us to benefit from the Gaussian splatting rendering for the scale-aware training by cross-view constraint and faster rendering on voxel grids as illustrated in Figure 2.
3.2 Scale-ware training by Gaussian Splatting
@@
Scale from spatial camera rig: Similar to the previous work [15, 44, 27], the scale information is from the surround camera rig. Specifically, the real-world scale can be obtained by leveraging camera extrinsic matrices which is to use spatial photometric loss in the overlap region between two adjacent views, i.e., warping ${I}^{i}_{t}$ to ${I}^{j}_{t}$:
$p^{i\rightarrow j}_{t}=K^{j}(T^{j})^{-1}T^{i}D_{t}^{i}(K^{i})^{-1}p^{i}_{t}$ (1)
@@
where $K^{i},T^{i}$ are the intrinsic and extrinsic matrices of $i$-th camera, $D_{t}^{i}$ is the predicted depth map of $i$-th camera, $p_{t}$ is the corresponding pixel during the warping. The warping operation is achieved by bilinear interpolation with the corresponding $p^{i\rightarrow j}_{t}$. However, as pointed out by [44], the mapping $p^{i\rightarrow j}_{t}$ can easily go out of image bounds with the weak supervision signal in such a small overlap region. Therefore, apart from the spatial loss, [44] proposes to facilitate the Structure-from-Motion (SFM) to extract sparse depth information for the direct depth supervision to provide a stronger supervision signal, but it is time-consuming and not straightforward. Different from [44], [27] enhances the depth estimation performance with spatio-temporal context that does not need the sparse depth from SFM but the performance is still limited. Inspired by the explicit sparse depth supervision in [44], we ask whether we can enforce the cross-view constraint on adjacent views more explicitly. The answer is yes. We find that the nature of Gaussian splatting is scale-aware projection that could serve for the cross-view stereo constraint. We propose Gaussian splatting for projection in stage 1 for better scale-aware training.
@@
Gaussian Splatting for Projection (GSP): As illustrated in Figure 2, we adapt a depth network [52] to predict the Gaussian attributes in 2D grid space, where, apart from the original depth map, we also predict the scale map and rotation map. For each adjacent view, we first calculate the mask in the overlap region, then mask out one side of these overlap regions. Due to the presence of the other side’s overlap region, the unprojected 3D scene remains complete, which is critical for providing scale training, as indicated in the experiment section (Table 3). We then perform splatting rendering on the adjacent views to obtain the rendered image. If the depth map is accurately learned, the rendered image should resemble the original images, providing the necessary scale information for the entire pipeline.
@@
The process of acquiring the overlap mask is illustrated in Figure 3. The acquisition of the overlap mask is similar to volume rendering. We densely sample points along the ray in one view, and a pixel is considered part of the overlap region if more than one sampled 3D point falls within the adjacent view. The overlap mask is determined by the camera extrinsic and defined max depth (e.g., 80m in nuScenes). Note that in the DDAD dataset [13], we exclude regions with self-occlusion (such as parts of the vehicle body). Finally, we apply an erosion operation from OpenCV library [2] to the mask for purification.
@@ 3.3 Fast rendering by Gaussian Splatting
Inefficient performance in volume rendering: For 2D supervision (semantic and depth maps), previous methods [11, 54, 23, 37] employed volume rendering based on dense sampling.Although the final 3D voxel representation for modeling the 3D scene is much quicker than the original implicit representation [35], it remains time-consuming, particularly when incorporating semantic map rendering. For example, in OccNeRF [54], the number of sampled points at a resolution of $180\times 320$ is 108,735,066. However, the target optimized points correspond to the vertices in the 3D voxel grid, which total $320\times 320\times 24=2,577,525$. This redundancy in densely sampled points helps optimization with volume rendering but is highly inefficient.
Gaussian Splatting from Voxel (GSV): As analyzed above, the target optimized points are the vertices in the 3D voxel grid, prompting us to consider directly optimizing these vertices. Interestingly, we find that it is suitable to use the Gaussian splatting to replace the volume rendering if we regard the vertices of each the voxel grid as the position of the 3D Gaussian. Then, we can optimize the attributes of the 3D Gaussian, such as semantic and opacity information. For example, in the empty space of the voxel, even though we have the vertices at that region during the splatting rendering, after the optimization, the network would predict the opacity as zero at these vertices then these vertices would not contribute any geometry or semantic information during the rendering. Since all the vertices are arranged in 3D voxel space with real-world 3D position $\mathcal{X}$, we use fixed scale $s$ and fixed rotation $r$ for each vertex for simplification. This allows us to model the 3D scene by optimizing the rest Gaussian attributes (semantic and opacity).
@@ 3.4 Loss function
$\mathcal{L}_{stage2}=\mathcal{L}_{temporal}+\mathcal{L}_{semantic},$ (3)
$(\mathcal{L}_{temporal},\mathcal{L}_{cross})=\frac{1-\operatorname{SSIM}\left(I_{t},\hat{I}_{t}\right)}{2}+(\beta)\left\|I_{t}-\hat{I}_{t}\right\|,$ (4)
where $I_{t}$ and $\hat{I}_{t}$ refer to the target image and the corresponding synthesized image, respectively. Note that $\hat{I}_{t}$ in the temporal-view photometric loss $\mathcal{L}_{temporal}$ is generated by projecting pixels from the source image using the coordinate index. In contrast, $\hat{I}_{t}$ in the cross-view photometric loss $\mathcal{L}_{cross}$ is derived from our proposed cross-view Gaussian splatting method. $\mathcal{L}_{semantic}$ is 2D semantic loss. The $\beta$ is set to 0.15 to balance the two loss terms. For $\mathcal{L}_{temporal}$ and $\mathcal{L}_{semantic}$, we recommend [54] for further details.
@@
Table 1: 3D occupancy prediction performance on the Occ3D-nuScenes dataset in mIoU metric. Since ‘other’ and ‘other flat’ classes are the invalid prompts for open-vocabulary models, we also calculate ‘mIoU*’ as the result ignoring the classes that do not consider these two classes during evaluation, while ‘mIoU’ is the original result. GT Occ. means using the ground truth occupancy label for the supervision. GT Pose is the ground truth pose from the sensor for self-supervised geometry learning. The full table is placed in the supplementary material with each category result.
@@
4 Experiment
4.1 Tasks, datasets, and metric
nuScenes [3]: For 3D occupancy estimation, we utilize annotations from Occ3D-nuScenes [41], which includes 600 scenes for training and 150 scenes for evaluation. Each sample has a perception range of [-40m, -40m, -1m, 40m, 40m, 5.4m] with a voxel size of 0.4m. For fair comparison, we use the 2D semantic map provided by OccNeRF [54]. We measure 3D occupancy estimation performance using the mean Intersection over Union (mIoU) metric. For depth estimation, we set the perception range in [-80m, -80m, -1m, 80m, 80m, 6m], while we clamp the ground truth to a range of 0.1m to 80m for evaluation, consistent with OccNeRF and SurroundDepth [44]. We evaluate depth maps using error metrics (Abs Rel, Sq Rel, RMSE, RMSE log) and threshold accuracy metrics ($\delta$).
DDAD [13]: Due to the lack of ground truth labels, we only present qualitative results for the fully self-supervised 3D occupancy estimation setting. Following the procedure on the nuScenes dataset, we obtain the 2D semantic labels according to the OccNeRF pipeline. For depth estimation, we also set the perception range in [-80m, -80m, -1m, 80m, 80m, 6m], but clamp the ground truth to a range of 0.1m to 200m for evaluation, consistent with SurroundDepth [44].
@@ 4.1 Tasks, datasets, and metric
Table 3: Ablation study for scale-aware depth estimation on the nuScenes dataset [14]. ✓* means the result from the original paper and ✓means the result using New-CRFs [52] as the depth network the same as ours. GS loss means using the spatial context constraint by our proposed Gaussian splatting for projection. Mask represents using the mask-out strategy before the unprojection. Erode means the erode process to the binary overlap mask and Refine is the refinement of depth estimation network with 2 epochs by fixing the 6D pose network.

### v4

@@ Abstract
We introduce GaussianOcc, a systematic method that investigates Gaussian splatting for fully self-supervised and efficient 3D occupancy estimation in surround views. First, traditional methods for self-supervised 3D occupancy estimation still require ground truth 6D ego pose from sensors during training. To address this limitation, we propose Gaussian Splatting for Projection (GSP) module to provide accurate scale information for fully self-supervised training from adjacent view projection. Additionally, existing methods rely on volume rendering for final 3D voxel representation learning using 2D signals (depth maps and semantic maps), which is time-consuming and less effective. We propose Gaussian Splatting from Voxel space (GSV) to leverage the fast rendering properties of Gaussian splatting. As a result, the proposed GaussianOcc method enables fully self-supervised (no ground truth ego pose) 3D occupancy estimation in competitive performance with low computational cost (2.7 times faster in training and 5 times faster in rendering). The relevant code is available in https://github.com/GANWANSHUI/GaussianOcc.git.
@@
In this supplementary material, we provide more implementation details, experiment results with analysis, and further discussion on the limitations and future work.
@@ 1 Introduction
Figure 1: Problem setting of GaussianOcc. Given a surround image sequence, the spatial camera extrinsic and its corresponding 2D semantic annotation, GaussianOcc is able to perform 3D occupancy estimation without the need for ground truth occupancy label and ground truth 6D ego pose for training.
Surround view 3D occupancy estimation [31, 45, 51, 48, 34, 39, 56] has emerged as a core perception task and a promising alternative to bird’s-eye view (BEV) methods [9, 28, 29]. To facilitate 3D occupancy estimation, several benchmarks have been developed for supervised training [45, 43, 41, 42], though these require substantial effort in 3D annotation. To reduce the burden of 3D annotation, self-supervised [11, 54, 22, 5, 16] and weakly-supervised [37] learning approaches based on volume rendering have been proposed [35, 46, 10]. Volume rendering allows 3D representation learning using 2D supervision signals, such as 2D semantic maps and depth maps, thereby eliminating the need for extensive 3D annotation.
Existing methods [54, 22] achieve self-supervised learning through volume rendering, where the 2D semantic map supervision is derived from open-vocabulary semantic segmentation [55], and the depth map supervision is obtained from self-supervised depth estimation [11, 12]. However, these approaches face two significant limitations. First, volume rendering is performed at real-world scale, which requires the ground truth 6D ego pose to calculate the multi-view photometric loss across sequential images. Second, the inefficiency in volume rendering poses a challenge, the same as in novel view synthesis tasks [26, 1, 17], due to the dense sampling operation required. These limitations impede the development of a more general and efficient paradigm for self-supervised 3D occupancy estimation.
To address the aforementioned limitations, we propose a fully self-supervised and efficient approach to 3D occupancy estimation with dedicated designs based on Gaussian splatting [26, 1]. Specifically, we introduce the use of Gaussian splatting to perform cross-view splatting, where the rendered image constructs a cross-view loss that provides scale information during joint training with the 6D pose network. This eliminates the need for ground truth 6D ego pose during training. To improve rendering efficiency, we move away from the dense sampling required in traditional volume rendering. Instead, we propose performing Gaussian splatting directly from the 3D voxel space. In this approach, each vertex in the voxel grid is treated as a 3D Gaussian, and we optimize the attributes of these Gaussians—such as semantic and opacity—directly within the voxel space. Through the above innovative approach, our proposed method makes progress toward fully self-supervised and efficient 3D occupancy estimation, as outlined in Figure 1.
@@
  • We propose Gaussian splatting for cross-view projection module, which can provide scale information to get rid of the need of ground truth 6D ego pose during the training.
@@ 2.2 Surround view 3D occupancy estimation
Figure 2: GaussianOcc is a two-stage method. In Stage 1, we train a scale-aware 6D pose network, using a U-Net architecture to predict Gaussian attributes in the 2D image grid space for cross-view Gaussian splatting. This approach provides scale information in the joint training with the 6D pose net. Based on the 6D pose from Stage 1, we perform self-supervised 3D occupancy estimation in Stage 2, where we lift the 2D features to a 3D voxel space and propose voxel grid Gaussian splatting for fast rendering. Note that, for clarity, we omit the line from the 6D pose network to the loss in Stage 1, and the 2D encoder is independent for each stage (not shared).
@@
Note that, recent works, GaussianFormer [23], GaussTR [24] and GaussianBeV [6], are related to ours in their focus on 3D occupancy estimation and BEV prediction. However, our exploration diverges by focusing on two new properties that Gaussian splatting can contribute to occupancy estimation: scale-aware training and faster rendering.
@@ 3.1 Preliminaries
3D Gaussian splatting [26] is for modeling static 3D scenes using point primitives, where each primitive is characterized with the following attributes: (1) a 3D position $\mathcal{X}\in\mathbb{R}^{3}$, (2) a color defined by SH coefficients $c\in\mathbb{R}^{k}$ (where $k$ denotes the dimensionality of the SH basis), (3) a rotation represented by a quaternion $r\in\mathbb{R}^{4}$, (4) a scaling factor $s\in\mathbb{R}^{3}_{+}$, and (5) an opacity $\alpha\in[0,1]$. The original Gaussian splatting is for scene-specific, fast 3D novel view synthesis, where the attributes of Gaussian points are optimized by the multi-view constraint. Differently, we study fully self-supervised and efficient 3D occupancy estimation by exploring the Gaussian attributes that are well-aligned in both 2D and 3D grids. Our design allows us to benefit from the Gaussian splatting rendering for the scale-aware training by cross-view constraint and faster rendering on voxel grids, as illustrated in Figure 2.
3.2 Scale-aware training by Gaussian Splatting
@@
Scale from spatial camera rig: Similar to the previous work [14, 44, 27], the scale information is from the surround camera rig. Specifically, the real-world scale can be obtained by leveraging camera extrinsic matrices, which is to use spatial photometric loss in the overlap region between two adjacent views, i.e., warping ${I}^{i}_{t}$ to ${I}^{j}_{t}$:
$p^{i\rightarrow j}_{t}=K^{j}(T^{j})^{-1}T^{i}D_{t}^{i}(K^{i})^{-1}p^{i}_{t},$ (1)
@@
where $K^{i},T^{i}$ are the intrinsic and extrinsic matrices of $i$-th camera, $D_{t}^{i}$ is the predicted depth map of $i$-th camera, $p_{t}$ is the corresponding pixel during the warping. The warping operation is achieved by direct bilinear interpolation with the corresponding $p^{i\rightarrow j}_{t}$. However, as pointed out by [44], the mapping in such a small overlap region, $p^{i\rightarrow j}_{t}$, can easily go into sub-optimal depth result, where we verify this in the experiment section in Figure 5. Therefore, apart from the spatial loss, [44] proposes to facilitate the Structure-from-Motion (SFM) to extract sparse depth information for direct depth supervision to provide a stronger supervision signal, but it is time-consuming and not straightforward. Different from [44], [27] enhances the depth estimation performance with spatio-temporal context that does not need the sparse depth from SFM but the performance is still limited. Inspired by the explicit sparse depth supervision in [44], we ask whether we can enforce the cross-view constraint on adjacent views more explicitly. The answer is yes. We find that the nature of Gaussian splatting is scale-aware projection that could serve for the cross-view stereo constraint. We propose Gaussian splatting for projection in stage 1 for better scale-aware training as follows.
@@
Gaussian Splatting for Projection (GSP): As illustrated in Figure 2, we adapt a depth network [52] to predict the Gaussian attributes in 2D grid space, where, apart from the original depth map, we also predict the scale map and rotation map. For each adjacent view, we first calculate the mask in the overlap region, then mask out one side of these overlap regions. Due to the presence of the other side’s overlap region, the unprojected 3D scene remains complete if the depth map is predicted well. This mask-out step is critical for providing scale training, as indicated in the experiment section (Table 3). We then perform splatting rendering on the adjacent views to obtain the rendered image. If the depth map is accurately learned, the rendered image should resemble the original images, providing the necessary scale information for the joint training with 6D pose net.
@@
Overlap Mask: The process of acquiring the overlap mask is illustrated in Figure 3. We densely sample points along the ray in one view, and a pixel is considered part of the overlap region if more than one sampled 3D point falls within the adjacent view. The overlap mask is only determined by the camera’s extrinsic and defined max depth (e.g., 80m in nuScenes). Note that in the DDAD dataset [13], we exclude regions with self-occlusion (such as parts of the vehicle body). Besides, we apply the erosion operation from OpenCV [2] to the mask for purification.
@@ 3.3 Fast rendering by Gaussian Splatting
Inefficient performance in volume rendering: For 2D supervision (semantic and depth maps), previous methods [11, 54, 22, 37] employed volume rendering based on dense sampling. Although the final 3D voxel representation for modeling the 3D scene is much quicker than the original implicit representation [35], it remains time-consuming, particularly when incorporating semantic map rendering. For example, in OccNeRF [54], the number of sampled points at a resolution of $180\times 320$ is 108,735,066. However, the target optimized points correspond to the vertices in the 3D voxel grid, totaling $300\times 300\times 24=2,160,000$. This redundancy in densely sampled points helps optimization with volume rendering but is highly inefficient.
Gaussian Splatting from Voxel (GSV): As analyzed above, the target optimized points are the vertices in the 3D voxel grid, prompting us to consider directly optimizing these vertices. Interestingly, we find that it is suitable to use the Gaussian splatting to replace the volume rendering if we regard the vertices of each the voxel grid as the position of the 3D Gaussian. Then, we can optimize the attributes of the 3D Gaussian, such as semantic and opacity information. For example, in the empty space of voxel, even though we have the vertices at that region during the splatting rendering, after the optimization, the network would predict the opacity as zero at these vertices then these vertices would not contribute any geometry or semantic information during the rendering. Since all vertices are arranged in 3D voxel space with real-world 3D position $\mathcal{X}$, we can use fixed scale $s$ and fixed rotation $r$ for each vertex for simplification. This allows us to model the 3D scene by optimizing the rest Gaussian attributes (semantic and opacity).
@@ 3.4 Loss function
$\mathcal{L}_{stage2}=\mathcal{L}_{temporal}+\lambda\mathcal{L}_{semantic},$ (3)
$(\mathcal{L}_{temporal},\mathcal{L}_{cross})=\frac{1-\operatorname{SSIM}\left(I_{t},\hat{I}_{t}\right)}{2}+\beta\left\|I_{t}-\hat{I}_{t}\right\|,$ (4)
where $I_{t}$ and $\hat{I}_{t}$ refer to the target image and the corresponding synthesized image, respectively. Note that $\hat{I}_{t}$ in the temporal-view photometric loss $\mathcal{L}_{temporal}$ is generated by projecting pixels from the source image using the coordinate index. In contrast, $\hat{I}_{t}$ in the cross-view photometric loss $\mathcal{L}_{cross}$ is derived from our proposed cross-view Gaussian splatting method. $\mathcal{L}_{semantic}$ is 2D semantic loss with balanced weight $\lambda=0.02$. For $\mathcal{L}_{cross}$ and $\mathcal{L}_{temporal}$, we set $\beta$ is set to 0.15 for weight balance the same as [54].
@@
4 Experiment
4.1 Tasks, datasets, and metric
nuScenes [3]: For 3D occupancy estimation, we utilize annotations from Occ3D [41]. For fair comparison, we use the 2D pseudo semantic map provided by OccNeRF [54] for training. We measure 3D occupancy estimation performance using the mean Intersection over Union (mIoU) metric. For depth estimation, we set the perception range in [-80m, -80m, -1m, 80m, 80m, 6m], while we clamp the ground truth to a range of 0.1m to 80m for evaluation, consistent with OccNeRF and SurroundDepth [44]. We evaluate depth maps using error metrics (Abs Rel, Sq Rel, RMSE, RMSE log) and threshold accuracy metrics ($\delta$).
DDAD [13]: Though we do not have the 3D occupancy labels on DDAD dataset, we can present qualitative results thanks to our fully self-supervised 3D occupancy estimation setting. We obtain the 2D pseudo semantic labels for training following OccNeRF pipeline in nuScenes dataset. For depth estimation, we clamp the depth range within 0.1m and 200m for evaluation, consistent with SurroundDepth [44].
@@ 4.1 Tasks, datasets, and metric
Table 1: 3D occupancy comparison on the Occ3D dataset with mIoU metric. Since ‘other’ and ‘other flat’ classes are the invalid prompts for open-vocabulary models, we also calculate ‘mIoU*’ as the result ignoring the classes that do not consider these two classes during evaluation, while ‘mIoU’ is the original result. GT Occ. means using the ground truth occupancy label for supervision. GT Pose is the ground truth 6D ego pose from the sensor for self-supervised geometry learning.
@@
Table 3: Ablation study for scale-aware depth estimation on the nuScenes dataset [3]. ✓* means the result from the original paper and ✓means the result using New-CRFs [52] as the depth network the same as ours. GS loss means using the spatial context constraint by our proposed Gaussian splatting for projection. Mask represents using the mask-out strategy before the unprojection. Erode means the erode process to the binary overlap mask and Refine is the refinement of depth estimation network with 2 epochs by fixing the 6D pose net.